OpenStudy (anonymous):

Medal and fan!! Solve please. 2x + 6 ≥ –4 A. x ≤ –8 B. x ≤ –5 C. x ≥ –5 D. x ≥ –8

OpenStudy (anonymous):

imagine ur inequality sign as = so it is easy to solve 2x+6=-4

OpenStudy (anonymous):

now solve for x

OpenStudy (anonymous):

B

OpenStudy (alprincenofl):

2x + 6 ≥ –4 2x ≥ -4 -6 2x ≥ -10 2x/2 ≥ -10/2 x ≥ -5 So the answer is "B"

OpenStudy (anonymous):

If one has access to the Mathematica then the Reduce fumction is just the ticket. \[\text{Reduce}[2 x+6\geq -4]\to x\geq -5 \]

OpenStudy (anonymous):

so its c

OpenStudy (anonymous):

@cooolkilla02 yeah it is c

OpenStudy (anonymous):

because ur are not dividing by a negative so ur inequality stays the same

OpenStudy (anonymous):

Josh has been running for 6 min and completes a lap in 7 min. Gabe has been running for 2 min and completes a lap in 9 min. How many more laps must Josh and Gabe complete to have run the same amount of time?